OpenStudy (anonymous):

Another Trigonometry question Help! Solve the equation in the interval [0°,360°). 4 sin^2θ=3 Please show your steps thank you :)

myininaya (myininaya):

Ok so first step is to divide 4 on both sides: \[\sin^2(\theta)=\frac{3}{4}\] Second step is to get rid of that square on the left side so take square root of both sides: \[\sin(\theta)=\pm \sqrt{\frac{3}{4}}\]

OpenStudy (anonymous):

4 sin^2θ=3 sin(t)^2=3/4 divide by 4 sin(t)=sqrt (3/4) take sqrt of both sides t=arcsin(sqrt (3/4)) take arc sin

myininaya (myininaya):

\[\sin(\theta)=\frac{\sqrt{3}}{2} \text{ or } \sin(\theta)=-\frac{\sqrt{3}}{2}\] Use unit circle :)

OpenStudy (anonymous):

Hmm..i am actually confused on what the questions is asking for....i understand your guys work, i am just lost on what the question is asking

myininaya (myininaya):

It wants you to find all values of theta that could satisfy the equation in the interval given

OpenStudy (anonymous):

ohhh, cause so when i use my unit circle i get... sin 30°, sin 150°, sin 210° and sin 330° ....is this right?

myininaya (myininaya):

\[\sin(30) \neq \frac{\sqrt{3}}{2}\] \[\sin(150) \neq \frac{\sqrt{3}}{2}\] \[\sin(210) \neq \frac{-\sqrt{3}}{2}\] \[\sin(330) \neq \frac{-\sqrt{3}}{2}\]

myininaya (myininaya):

\[\sin(60)=?\] \[\sin(120)=?\] \[\sin(240)=?\] \[\sin(300)=?\]

OpenStudy (anonymous):

now I understand...okay i was looking at the ones in the x coordinate instead of the y coordinate...got it now

OpenStudy (anonymous):

what about sin 45 that would work also right?

myininaya (myininaya):

no \[\sin(45)=\frac{\sqrt{2}}{2} \neq \frac{\sqrt{3}}{2} \text{ or } \frac{-\sqrt{3}}{2}\]

myininaya (myininaya):

you are looking for where you have that the sine value is either \[\frac{\sqrt{3}}{2} \text{ or } \frac{-\sqrt{3}}{2}\]

OpenStudy (anonymous):

I see....now I understand, thanks again!!